OpenStudy (anonymous):

Find the power series of the function:

OpenStudy (anonymous):

\[f(x)=\frac{ x ^{3} }{ (x-9)^{2} }\]

OpenStudy (anonymous):

ok i'd probably say use this power series\[\frac{1}{1-u}=1+u+u^2+...\]for \(|u|<1\)

OpenStudy (anonymous):

I know to use geometric, I just don't know how to turn it into that form.

OpenStudy (anonymous):

start with\[\frac{1}{1-\frac{x}{9}}=1+\frac{x}{9}+\frac{x^2}{81}+...\]and then differentiate both sides

OpenStudy (anonymous):

for \(|x|<9\)

OpenStudy (anonymous):

Why do I start with that?

OpenStudy (anonymous):

I understand that you're changing 1/(9-x) to geometric form

OpenStudy (anonymous):

because its what u have in denum of first expression\[\frac{1}{1-\frac{x}{9}}\times (-\frac{1}{9})=\frac{1}{x-9}\]

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

But wouldn't you have to multiply it to the top and bottom?

OpenStudy (anonymous):

\[\frac{ 1 }{ 1-\frac{ x }{ 9 } }\times \frac{ -1/9 }{ -1/9 }\]

OpenStudy (anonymous):

i just wanted to say it can be change to what u have...yes actually u should multiply top and bot

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So you work from the geometric and then transition it to what you have? You don't try to manipulate the original equation?

OpenStudy (anonymous):

I thought we have to somehow manipulate it via deriving or integrating.

OpenStudy (anonymous):

thats what came in to my mind first...maybe there is a better way :)

OpenStudy (anonymous):

I've been so stuck on this problem for over an hour. :/

OpenStudy (anonymous):

ok i hope that helps :) good luck